didnt knnow where to place this problem...sorry
What does it mean to refer to a 20 in tv set or a 25 in tv set? Such units refer to the diagonal of the screen. A 30 in tv set also has a width of 24 inches. What is its height?
What is the height of a 30 in tv? ????inches

Answer by nerdybill(5412)

The measurement of a TV is usually done diagonally. Therefore, the width, length, and the diagonal forms a "right triangle". You can now apply Pythagorean theorem to solve your problem.
.
A 30 in tv set also has a width of 24 inches. What is its height?
30^2 = 24^2 + h^2
30^2 = 24^2 + h^2
900 = 576 + h^2
324 = h^2
18 inches = h (height of TV)